OpenStudy (anonymous):

find all the differential equations of all tangents to y=x^2

OpenStudy (anonymous):

The tangent at \(x_0\) is :\[ y-f(x_0) = f'(x_0) (x-x_0) \implies y = xf'(x_0)-x_0f'(x_0)+f(x_0) \]

OpenStudy (anonymous):

First find the derivative.

OpenStudy (anonymous):

after that?

OpenStudy (anonymous):

Plug in \(x_0\)

OpenStudy (anonymous):

done

OpenStudy (anonymous):

What is the final answer?

OpenStudy (anonymous):

but wherz the differential equation ?>

OpenStudy (anonymous):

y=x^2 applying d/dx on both sides \[\frac{ d }{ dx }y=\frac{ d }{ dx }x^2\] \[y \prime=2\times x ^{2-1}\] \[y \prime=2x\] \[\frac{ d }{ dx }y=2x\] remember this formula \[ \frac{d }{ dx }x ^{n}=nx ^{n-1}\]